OpenStudy (anonymous):

What is the diameter of a circle with area equal to one-hundred sixty nine pi meters squared?

OpenStudy (anonymous):

\[\large A = 169 \pi \Rightarrow 169\pi = \pi \times (radius)^2 \Rightarrow 169=(radius)^2 \Rightarrow 13 = radius \]

OpenStudy (anonymous):

can you find the diameter from here?